COMPUTER RECOGNITION OF THREE-MANIFOLDS

SERGEI V. MATVEEV

Abstract. We describe the theoretical background for a computer program that recognizes all closed orientable three-manifolds up to complexity 8. The program can treat also nonclosed threemanifolds and manifolds of greater complexity, but without necessarily succeeding in recognizing them.